# The universal multiplicity function: counting halos and voids

### v1

@@ Appendix A Standard deviation of the derivative of the filtered density contrast field
$\langle\delta^{\prime 2}_{S}\rangle=\left(\frac{{\rm d}R}{{\rm d}S}\right)^{2}\left\langle\left(\frac{\partial\delta(R)}{\partial R}\right)^{2}\right\rangle,$ (A1)
where, considering Eqs. (1)–(2),
$\left\langle\left(\frac{\partial\delta(R)}{\partial R}\right)^{2}\right\rangle=\int_{0}^{\infty}\frac{\mathrm{d}k\,k^{2}}{2\pi^{2}}P(k)\left(\frac{\partial W(kR)}{\partial R}\right)^{2}.$ (A2)
For the top-hat filter function, the corresponding derivative is
$\frac{\partial W(kR)}{\partial R}=\frac{3}{Rx^{3}}\left[\left(x^{2}-3\right)\sin(x)-3x\cos(x)\right]\,,$ (A3)
with $x=kR$. Note that for $k\gg 1/R$, this function behaves as $\sim\sin(kR)/k$. Therefore, for large $k$, the integrand of Eq. (A2) behaves as $\sim P(k)\sin^{2}(kR)$. As a result, the convergence of the integral relies solely on the behavior of $P(k)$ at large $k$, without any additional suppression term; therefore, in the cases where the integral convergences, the convergence is slow (see Appendix B in Paranjape et al., 2012). This makes the direct evaluation of this integral computationally expensive, as it requires the knowledge of $P(k)$ up to very high $k$ values. As a consequence, in the literature, this quantity is generally not computed with a top-hat filter; even when the top-hat filter is used for other quantities, such as $\langle\delta^{2}(R)\rangle$. Usually, $\langle\delta^{\prime 2}_{S}\rangle$ is evaluated with a Gaussian filter smoothed at $R_{\rm G}$, and mapped to a different top-hat scale $R_{\rm TH}$ according to various prescriptions (Porciani et al., 1998; Paranjape et al., 2013b, a; Biagetti et al., 2014).
@@
$\begin{split}\left\langle\frac{\partial\delta(R_{1})}{\partial R_{1}}\frac{\partial\delta(R_{2})}{\partial R_{2}}\right\rangle&=\frac{\partial}{\partial R_{1}}\frac{\partial}{\partial R_{2}}\left\langle\delta(R_{1})\delta(R_{2})\right\rangle\\ &=\frac{\partial}{\partial R_{1}}\frac{\partial}{\partial R_{2}}C(R_{1},R_{2})\,,\end{split}$ (A4)
where $C(R_{1},R_{2})$ is the covariance defined in Eq. (4), which is a well defined quantity. From that, the variance of the scale derivative of the density contrast field can be easily derived as
$\left\langle\left(\frac{\partial\delta(R)}{\partial R}\right)^{2}\right\rangle=\left.\frac{\partial}{\partial R}\frac{\partial}{\partial R^{\prime}}C(R,R^{\prime})\right|_{R=R^{\prime}}.$ (A5)
@@
$\left\langle\frac{\partial^{n}\delta(R)}{\partial R^{n}}\frac{\partial^{m}\delta(R)}{\partial R^{m}}\right\rangle=\left.\frac{\partial^{n}}{\partial R^{n}}\frac{\partial^{m}}{\partial R^{\prime m}}C(R,R^{\prime})\right|_{R=R^{\prime}}.$ (A6)
To further confirm the above relation, we measure the ensemble variance over many Monte Carlo realizations of the field $\partial\delta(R)/\partial R$, obtained by finite differentiation of the field $\delta(R)$ from Eq. (5) and using the top-hat filter. Fig. 7 shows the results: $\langle(\partial\delta(R)/\partial R)^{2}\rangle$ computed as the ensemble variance of Monte Carlo realizations (orange line); same quantity computed with Eq. (A5) (black dashed line); $\sigma^{2}(R)=\langle\delta^{2}(R)\rangle$ (blue line).
Figure 7: $\langle(\partial\delta(R)/\partial R)^{2}\rangle$ computed as the ensemble variance of Monte Carlo realizations (orange line); same quantity computed with Eq. (A5) (black dashed line); $\sigma^{2}(R)=\langle\delta^{2}(R)\rangle$ (blue line). All the quantity showed are computed with a top-hat filter
@@ Appendix B Derivation of the analytical multiplicity function
$\begin{split}p(&\delta_{S},\delta^{\prime}_{S},\delta_{s})=\\[-12.91663pt] &\frac{1}{(2\pi)^{3/2}\sqrt{\det{\bf\Sigma}}}\exp\left\{-\frac{1}{2}(\delta_{S},\delta^{\prime}_{S},\delta_{s})\begin{array}[]{c}\text{\LARGE$\bf\Sigma$}^{-1}\end{array}\left(\begin{array}[]{c}\delta_{S}\\ \delta^{\prime}_{S}\\ \delta_{s}\end{array}\right)\right\}\end{split}$ (B1)
where ${\bf\Sigma}$ is the covariance matrix of Eq. (10). The exponent of Eq. (B1) reads
$\frac{-1}{2\det{\bf\Sigma}}\Big{[}$ $\Gamma_{BB}B_{S}^{2}+\Gamma_{\delta^{\prime}\delta^{\prime}}\delta^{\prime 2}_{S}+\Gamma_{\delta\delta}\delta^{2}_{s}$ (B2)
$+2\Gamma_{B\delta^{\prime}}B_{S}\delta^{\prime}_{S}+2\Gamma_{B\delta}B_{S}\delta_{s}+2\Gamma_{\delta\delta^{\prime}}\delta^{\prime}_{S}\delta_{s}\Big{]}$
@@
$\det{\bf\Sigma}=sSD-SC^{\prime 2}+CC^{\prime}-s/4-DC^{2}$ (B3)
and $\Sigma_{ij}^{-1}=\Gamma_{ij}/\det{\bf\Sigma}$, with $\Gamma_{ij}$ are listed in Eq. (12). The integral with respect $\delta^{\prime}$ of Eq. (9) is
$\int_{B^{\prime}}^{\infty}$ ${\rm d}\delta^{\prime}p(B_{S},\delta^{\prime}_{S},\delta_{s})(\delta^{\prime}_{S}-B^{\prime}_{S})=\frac{1}{(2\pi)^{3/2}\sqrt{\det{\bf\Sigma}}}\times$
@@
$\sqrt{\frac{\pi}{2}\frac{\det{\bf\Sigma}}{\Gamma_{\delta^{\prime}\delta^{\prime}}}}\left(B^{\prime}_{S}+\frac{\kappa}{\Gamma_{\delta^{\prime}\delta^{\prime}}}\right)\exp\left[\frac{\kappa^{2}/\Gamma_{\delta^{\prime}\delta^{\prime}}}{2\det{\bf\Sigma}}\right]\times$
@@
we obtain Eq. (3).
Appendix C Global and boundary behavior
$\boldsymbol{s\rightarrow S}$ limit
Eq. (3) behaves as $/0$ 0 as $s\rightarrow S$. The quantities $\Gamma_{ij}$ and $\det{\bf\Sigma}$ can be written in terms of $s$, $S$, $D$, $C$, $C^{\prime}$. In particular, $C$ and $C^{\prime}$ can be expanded around $S$ as
$\begin{split}C(s)&=\sum_{n}\frac{(s-S)^{n}}{n!}\langle\delta\delta^{(n)}\rangle|_{s=S}\\ C^{\prime}(s)&=\sum_{n}\frac{(s-S)^{n}}{n!}\langle\delta^{\prime}\delta^{(n)}\rangle|_{s=S},\end{split}$ (C1)
@@ Appendix C Global and boundary behavior
  • $\langle\delta\delta^{\prime\prime}\rangle$: $\frac{\partial}{\partial S}\langle\delta\delta^{\prime}\rangle=0=\langle\delta\delta^{\prime\prime}\rangle+\langle\delta^{\prime 2}\rangle$
           $\Rightarrow\langle\delta\delta^{\prime\prime}\rangle=\langle\delta^{\prime 2}\rangle=-D$
@@
  • $\langle\delta\delta^{\prime\prime\prime}\rangle$: $\frac{\partial^{2}}{\partial S^{2}}\langle\delta\delta^{\prime}\rangle=0=\frac{\partial}{\partial S}\left[\langle\delta\delta^{\prime\prime}\rangle+\langle\delta^{\prime 2}\rangle\right]$
             $=\langle\delta\delta^{\prime\prime\prime}\rangle+3\langle\delta^{\prime}\delta^{\prime\prime}\rangle$
           $\Rightarrow\langle\delta\delta^{\prime\prime\prime}\rangle=-\frac{3}{2}D^{\prime}$
  • $\langle\delta^{\prime}\delta^{\prime\prime\prime}\rangle$: $\frac{\partial}{\partial S}\langle\delta^{\prime}\delta^{\prime\prime\prime}\rangle=\langle\delta^{\prime}\delta^{\prime\prime\prime}\rangle+\langle\delta^{\prime\prime 2}\rangle$
           $\Rightarrow\langle\delta^{\prime}\delta^{\prime\prime\prime}\rangle=\frac{1}{2}D^{\prime\prime}-\langle\delta^{\prime\prime 2}\rangle$
@@
  • $\langle\delta\delta^{\prime\prime\prime\prime}\rangle$: $\frac{\partial}{\partial S}\left[\langle\delta\delta^{\prime\prime\prime}\rangle+3\langle\delta^{\prime}\delta^{\prime\prime}\rangle\right]=0$
              $=\langle\delta\delta^{\prime\prime\prime\prime}\rangle+4\langle\delta^{\prime}\delta^{\prime\prime\prime}\rangle+3\langle\delta^{\prime\prime 2}\rangle$
           $\Rightarrow\langle\delta\delta^{\prime\prime\prime\prime}\rangle=\langle\delta^{\prime\prime 2}\rangle+2D^{\prime\prime}$
  • $\langle\delta^{\prime}\delta^{\prime\prime\prime\prime}\rangle$: $\frac{\partial^{3}}{\partial S^{3}}\langle\delta^{\prime 2}\rangle=6\langle\delta^{\prime\prime}\delta^{\prime\prime\prime}\rangle+2\langle\delta^{\prime}\delta^{\prime\prime\prime\prime}\rangle$
           $\Rightarrow\langle\delta^{\prime}\delta^{\prime\prime\prime\prime}\rangle=\frac{1}{2}D^{\prime\prime\prime}-\frac{3}{2}\frac{\partial}{\partial S}\langle\delta^{\prime\prime 2}\rangle$
@@
where $\langle\delta^{\prime\prime 2}\rangle$ can be evaluated following Appendix A. Using these relations, it is possible to Fourier expand $\Gamma_{ij}$, Eq. (12), and ${\det{\bf\Sigma}}$, Eq. (10), in the limit $s\rightarrow S$. The first non-zero coefficient of the Taylor expansion around $S$ of ${\det{\bf\Sigma}}$ is the $4^{\rm th}$ one, i.e. ${\det{\bf\Sigma}}\propto(s-S)^{4}$ as $s\rightarrow S$.
For convenience, we rewrite Eq. (3) as
${\cal P}(s)=I_{1}+I_{2}+e^{-B^{2}_{S}/2S}\left(I_{3}+I_{4}\right),$ (C2)
@@
where the 4 terms $I_{i}$ are in the same order of the ones appearing in the original equation. $I_{1}$, the first term on the right-hand-side of Eq. (3), i.e. the second and third lines, goes to zeros, as the exponential and $\Gamma_{\delta\delta}$ are constant, while $\Gamma_{\delta^{\prime}\delta^{\prime}}\propto(s-S)^{2}$. Therefore, the quantity ${\det{\bf\Sigma}}/\Gamma_{\delta^{\prime}\delta^{\prime}}\propto(s-S)^{2}$, which multiplies a finite quantity, i.e. $I_{1}\rightarrow 0$ as $s\rightarrow S$. Expanding up to the second order, the second term, $I_{2}$, i.e. the fourth line of Eq. (3), becomes
@@
$I_{2}\rightarrow\frac{\sqrt{\Gamma_{\delta\delta}}}{4\pi S}\exp\left(-\frac{DB_{S}^{2}+SB^{\prime}-B_{S}B}{2\Gamma_{\delta\delta}}\right).$ (C3)
@@
$\sqrt{\frac{S}{2\Gamma_{\delta^{\prime}\delta^{\prime}}}}\left(B(s)-\frac{C}{S}B_{S}\right)\rightarrow\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\left(\frac{B_{S}}{2S}-B^{\prime}\right).$ (C4)
@@
$I_{3}=\frac{\det{\bf\Sigma}}{2\pi\sqrt{\Gamma_{\delta^{\prime}\delta^{\prime}}}\Gamma_{\delta^{\prime}\delta}}{\cal F}(x,\alpha,\beta)$ (C5)
where ${\det{\bf\Sigma}}/(\sqrt{\Gamma_{\delta^{\prime}\delta^{\prime}}}\Gamma_{\delta^{\prime}\delta})\propto(s-S)^{2}$, and ${\cal F}(x,\alpha,\beta)$ is defined in Eq. (13). It can be noticed from Eq. (14) that
$x\rightarrow\sqrt{\frac{\Gamma_{\delta\delta}}{2\det{\bf\Sigma}_{(4)}}}$ $\left[\left(\frac{D^{\prime}}{2}-D^{2}\right)\frac{B_{s}}{2\Gamma_{\delta\delta}}\right.-$ (C6)
@@
where $\det{\bf\Sigma}_{(4)}$ is the $4^{\rm th}$-order coefficient of the Taylor expansion. The above equation is a finite quantity, while $\alpha\propto(s-S)^{2}$ and $\beta\propto(s-S)^{-1}$. It follows that in the limit $s\rightarrow S$, the exponential term appearing in the integral of Eq. (13) behaves as a Gaussian infinitely wide and centered at $\infty$- ∞. Note, however, that the width of this Gaussian, i.e. $\sqrt{2/\alpha}$, and the center position, i.e. $-\beta$, compensate for each other,
$\alpha\beta^{2}=S\frac{\Gamma_{\delta^{\prime}\delta^{\prime}}}{\Gamma_{\delta^{\prime}\delta}^{2}}\left(\frac{B_{S}}{2S}-B^{\prime}\right)^{2}\rightarrow\frac{S}{\Gamma_{\delta\delta}}\left(\frac{B_{S}}{2S}-B^{\prime}\right)^{2}.$ (C7)
This means that the exponential contributes to the integral as a finite positive quantity in the entire range from $\infty$- ∞ to $x$. As a consequence, the upper integration limit, $x$, is negligible and can be considered $x\rightarrow 0$. The same argument applies to the error function: the scale range in which it differs from -1 is negligible, therefore we can consider ${\rm erf}(t)\rightarrow-1$ over all the scales of interest. In this limit, the integral of Eq. (13) becomes
${\cal F}(x,\alpha,\beta)$ $\rightarrow-\int_{-\infty}^{0}{\rm d}t\,t\,e^{-\alpha(t+\beta)^{2}}$ (C8)
@@
from which, the limit for $s\rightarrow S$ of Eq. (C5), is
$I_{3}\rightarrow\frac{\sqrt{\Gamma_{\delta\delta}}}{4\pi S}e^{-\alpha\beta^{2}}+\frac{{\rm erf}(\sqrt{\alpha}\beta)+1}{4\sqrt{s\pi S}}\left(\frac{B_{S}}{2S}-B^{\prime}\right)\,.$ (C9)
Substituting the four $I_{i}$ terms computed here in Eq. (C2), we find that Eq. (3), in the limit $s\rightarrow S$, becomes
$\lim_{s\rightarrow S}{\cal P}(s)=$ (C10)
$e^{-B^{2}_{S}/2S}\left\llbracket\frac{\sqrt{\Gamma_{\delta\delta}}}{2\pi S}\exp\left[-\frac{S}{2\Gamma_{\delta\delta}}\left(\frac{B_{S}}{2S}-B^{\prime}_{S}\right)^{2}\right]\right.+$
@@
$\frac{DB_{S}^{2}+SB^{\prime}-B_{S}B}{2\Gamma_{\delta\delta}}=\frac{S}{2\Gamma_{\delta\delta}}\left(\frac{B_{S}}{2S}-B^{\prime}_{S}\right)^{2}+\frac{B_{S}}{2S}.$ (C11)
$\boldsymbol{s\rightarrow 0}$ limit
We now explore the $s\rightarrow 0$ limit of ${\cal P}(s)$, Eq. (3). Note that $S>s$ by construction, so $R_{S}<R_{s}$. Considering $S$ as a finite quantity, $R_{S}\ll R_{s}\rightarrow\infty$ as $s\rightarrow 0$, entailing that
$C$ $=\int\frac{{\rm d}k}{2\pi^{2}}k^{2}P(k)W(kR_{S})W^{*}(kR_{s})$ (C12)
$\rightarrow\int\frac{{\rm d}k}{2\pi^{2}}k^{2}P(k)W(kR_{s})={\cal O}(s)\,,$ (C13)

### v2

@@ Appendix A Correlation of the density contrast field scale derivatives
$\langle\delta^{\prime 2}_{S}\rangle=\left(\frac{{\rm d}R}{{\rm d}S}\right)^{2}\left\langle\left(\frac{\partial\delta(R)}{\partial R}\right)^{2}\right\rangle,$ (A.1)
where, considering Eqs. (2.1)–(3.1),
$\left\langle\left(\frac{\partial\delta(R)}{\partial R}\right)^{2}\right\rangle=\int_{0}^{\infty}\frac{\mathrm{d}k\,k^{2}}{2\pi^{2}}P(k)\left(\frac{\partial W(kR)}{\partial R}\right)^{2}.$ (A.2)
The derivative of the top-hat filter function with respect to the smoothing lenght $R$ is
$\frac{\partial W(kR)}{\partial R}=\frac{3}{Rx^{3}}\left[\left(x^{2}-3\right)\sin(x)+3x\cos(x)\right]\,,$ (A.3)
with $x=kR$. Note that for $k\gg 1/R$, this function behaves as $\sim\sin(kR)/k$. Therefore, for large $k$, the integrand of Eq. (A.2) behaves as $\sim P(k)\sin^{2}(kR)$. As a result, the convergence of the integral relies solely on the behaviour of $P(k)$ at large $k$, without any additional suppression term; therefore, in the cases where the integral convergences, the convergence is slow [see Appendix B in 19]. This makes the direct evaluation of this integral computationally expensive, as it requires the knowledge of $P(k)$ up to very high $k$ values. As a consequence, in the literature, this quantity is generally not computed with a top-hat filter; even when the top-hat filter is used for other quantities, such as $\langle\delta^{2}(R)\rangle$. Usually, $\langle\delta^{\prime 2}_{S}\rangle$ is evaluated with a Gaussian filter smoothed at $R_{\rm G}$, and mapped to a different top-hat scale $R_{\rm TH}$ according to various prescriptions [7, 31, 93, 94].
@@
$\begin{split}\left\langle\frac{\partial\delta(R_{1})}{\partial R_{1}}\frac{\partial\delta(R_{2})}{\partial R_{2}}\right\rangle&=\frac{\partial}{\partial R_{1}}\frac{\partial}{\partial R_{2}}\left\langle\delta(R_{1})\delta(R_{2})\right\rangle\\ &=\frac{\partial}{\partial R_{1}}\frac{\partial}{\partial R_{2}}C(R_{1},R_{2})\,,\end{split}$ (A.4)
where $C(R_{1},R_{2})$ is the covariance defined in Eq. (3.3), which is a well defined quantity. From that, the variance of the scale derivative of the density contrast field can be easily derived as
$\left\langle\left(\frac{\partial\delta(R)}{\partial R}\right)^{2}\right\rangle=\left.\frac{\partial}{\partial R_{1}}\frac{\partial}{\partial R_{2}}C(R_{1},R_{2})\right|_{R_{1}=R_{2}=R}.$ (A.5)
@@
$\left\langle\frac{\partial^{n}\delta(R)}{\partial R^{n}}\frac{\partial^{m}\delta(R)}{\partial R^{m}}\right\rangle=\left.\frac{\partial^{n}}{\partial R_{1}^{n}}\frac{\partial^{m}}{\partial R_{2}^{m}}C(R_{1},R_{2})\right|_{R_{1}=R_{2}=R}.$ (A.6)
To further confirm the above relation, we measure the ensemble variance over many MC realisations of the field $\partial\delta(R)/\partial R$, obtained by finite differentiation of the field $\delta(R)$ from Eq. (3.4) with the top-hat filter. Fig. 7 shows the results: $\langle(\partial\delta(R)/\partial R)^{2}\rangle$ computed as the ensemble variance of MC realisations (orange line); same quantity computed with Eq. (A.5) (black dashed line); $\sigma^{2}(R)=\langle\delta^{2}(R)\rangle$ (blue line).
Figure 7: $\langle(\partial\delta(R)/\partial R)^{2}\rangle$ computed as the ensemble variance of MC realisations (orange line); same quantity computed with Eq. (A.5) (black dashed line); $\sigma^{2}(R)=\langle\delta^{2}(R)\rangle$ (blue line). All the quantity showed are computed with a top-hat filter
@@ Appendix B Derivation of the analytical multiplicity function
$p(\delta_{S},\delta^{\prime}_{S},\delta_{s})=\frac{1}{(2\pi)^{3/2}\sqrt{\det{\bf\Sigma}}}\exp\left\{-\frac{1}{2}(\delta_{S},\delta^{\prime}_{S},\delta_{s})\begin{array}[]{c}\text{\LARGE$\bf\Sigma$}^{-1}\end{array}\left(\begin{array}[]{c}\delta_{S}\\ \delta^{\prime}_{S}\\ \delta_{s}\end{array}\right)\right\}$ (B.1)
where ${\bf\Sigma}$ is the covariance matrix of Eq. (3.9). The exponent of Eq. (B.1) reads
$\frac{-1}{2\det{\bf\Sigma}}\left[\Gamma_{BB}B_{S}^{2}+\Gamma_{\delta^{\prime}\delta^{\prime}}\delta^{\prime 2}_{S}+\Gamma_{\delta\delta}\delta^{2}_{s}+2\Gamma_{B\delta^{\prime}}B_{S}\delta^{\prime}_{S}+2\Gamma_{B\delta}B_{S}\delta_{s}+2\Gamma_{\delta\delta^{\prime}}\delta^{\prime}_{S}\delta_{s}\right]$ (B.2)
@@
$\det{\bf\Sigma}=sSD-SC^{\prime 2}+CC^{\prime}-s/4-DC^{2}$ (B.3)
and $\Sigma_{ij}^{-1}=\Gamma_{ij}/\det{\bf\Sigma}$, with $\Gamma_{ij}$ listed in Eq. (3.11). The integral with respect $\delta^{\prime}$ of Eq. (3.8) is
$\int_{B^{\prime}}^{\infty}$ ${\rm d}\delta^{\prime}p(B_{S},\delta^{\prime}_{S},\delta_{s})(\delta^{\prime}_{S}-B^{\prime}_{S})=$ (B.4)
@@
we obtain Eq. (3.2).
Appendix C Global and boundary behaviour
C.1 $s\rightarrow S$ limit
Eq. (3.2) behaves as $/0$ 0 as $s\rightarrow S$. The quantities $\Gamma_{ij}$ and $\det{\bf\Sigma}$ can be written in terms of $s$, $S$, $D$, $C$, $C^{\prime}$. In particular, $C$ and $C^{\prime}$ can be expanded around $S$ as
$\begin{split}C(s)&=\sum_{n}\frac{(s-S)^{n}}{n!}\langle\delta\delta^{(n)}\rangle|_{s=S}\\ C^{\prime}(s)&=\sum_{n}\frac{(s-S)^{n}}{n!}\langle\delta^{\prime}\delta^{(n)}\rangle|_{s=S},\end{split}$ (C.1)
@@ C.1 $s\rightarrow S$ limit
  • $\langle\delta\delta^{\prime\prime}\rangle$: $\frac{\partial}{\partial S}\langle\delta\delta^{\prime}\rangle=0=\langle\delta\delta^{\prime\prime}\rangle+\langle\delta^{\prime 2}\rangle\Rightarrow\langle\delta\delta^{\prime\prime}\rangle=\langle\delta^{\prime 2}\rangle=-D$
@@
  • $\langle\delta\delta^{\prime\prime\prime}\rangle$: $\frac{\partial^{2}}{\partial S^{2}}\langle\delta\delta^{\prime}\rangle=0=\frac{\partial}{\partial S}\left[\langle\delta\delta^{\prime\prime}\rangle+\langle\delta^{\prime 2}\rangle\right]=\langle\delta\delta^{\prime\prime\prime}\rangle+3\langle\delta^{\prime}\delta^{\prime\prime}\rangle\Rightarrow\langle\delta\delta^{\prime\prime\prime}\rangle=-\frac{3}{2}D^{\prime}$
  • $\langle\delta^{\prime}\delta^{\prime\prime\prime}\rangle$: $\frac{\partial}{\partial S}\langle\delta^{\prime}\delta^{\prime\prime\prime}\rangle=\langle\delta^{\prime}\delta^{\prime\prime\prime}\rangle+\langle\delta^{\prime\prime 2}\rangle\Rightarrow\langle\delta^{\prime}\delta^{\prime\prime\prime}\rangle=\frac{1}{2}D^{\prime\prime}-\langle\delta^{\prime\prime 2}\rangle$
@@
  • $\langle\delta\delta^{\prime\prime\prime\prime}\rangle$: $\frac{\partial}{\partial S}\left[\langle\delta\delta^{\prime\prime\prime}\rangle+3\langle\delta^{\prime}\delta^{\prime\prime}\rangle\right]=0=\langle\delta\delta^{\prime\prime\prime\prime}\rangle+4\langle\delta^{\prime}\delta^{\prime\prime\prime}\rangle+3\langle\delta^{\prime\prime 2}\rangle\Rightarrow\langle\delta\delta^{\prime\prime\prime\prime}\rangle=\langle\delta^{\prime\prime 2}\rangle+2D^{\prime\prime}$
  • $\langle\delta^{\prime}\delta^{\prime\prime\prime\prime}\rangle$: $\frac{\partial^{3}}{\partial S^{3}}\langle\delta^{\prime 2}\rangle=6\langle\delta^{\prime\prime}\delta^{\prime\prime\prime}\rangle+2\langle\delta^{\prime}\delta^{\prime\prime\prime\prime}\rangle\Rightarrow\langle\delta^{\prime}\delta^{\prime\prime\prime\prime}\rangle=\frac{1}{2}D^{\prime\prime\prime}-\frac{3}{2}\frac{\partial}{\partial S}\langle\delta^{\prime\prime 2}\rangle$
@@
where $\langle\delta^{\prime\prime 2}\rangle$ can be evaluated following Appendix A. Using these relations, it is possible to Fourier expand $\Gamma_{ij}$, Eq. (3.11), and ${\det{\bf\Sigma}}$, Eq. (3.9), in the limit $s\rightarrow S$. The first non-zero coefficient of the Taylor expansion around $S$ of ${\det{\bf\Sigma}}$ is the $4^{\rm th}$ one, i.e. ${\det{\bf\Sigma}}\propto(s-S)^{4}$ as $s\rightarrow S$.
For convenience, we rewrite Eq. (3.2) as
${\cal P}(s)=I_{1}+I_{2}+e^{-B^{2}_{S}/2S}\left(I_{3}+I_{4}\right),$ (C.2)
@@
where the 4 terms $I_{i}$ are in the same order of the ones appearing in Eq. (3.2). $I_{1}$, the first term on the right-hand-side of Eq. (3.2), i.e. the second and third lines, goes to zeros, as the exponential and $\Gamma_{\delta\delta}$ are constant, while $\Gamma_{\delta^{\prime}\delta^{\prime}}\propto(s-S)^{2}$. Therefore, the quantity ${\det{\bf\Sigma}}/\Gamma_{\delta^{\prime}\delta^{\prime}}\propto(s-S)^{2}$, which multiplies a finite quantity, i.e. $I_{1}\rightarrow 0$ as $s\rightarrow S$. Expanding up to the second order, the second term, $I_{2}$, i.e. the fourth line of Eq. (3.2), becomes
@@
$I_{2}\rightarrow\frac{\sqrt{\Gamma_{\delta\delta}}}{4\pi S}\exp\left(-\frac{DB_{S}^{2}+SB^{\prime}-B_{S}B}{2\Gamma_{\delta\delta}}\right).$ (C.3)
@@
$\sqrt{\frac{S}{2\Gamma_{\delta^{\prime}\delta^{\prime}}}}\left(B(s)-\frac{C}{S}B_{S}\right)\rightarrow\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\left(\frac{B_{S}}{2S}-B^{\prime}\right).$ (C.4)
@@
$I_{3}=\frac{\det{\bf\Sigma}}{2\pi\sqrt{\Gamma_{\delta^{\prime}\delta^{\prime}}}\Gamma_{\delta^{\prime}\delta}}{\cal F}(x,\alpha,\beta)$ (C.5)
where ${\det{\bf\Sigma}}/(\sqrt{\Gamma_{\delta^{\prime}\delta^{\prime}}}\Gamma_{\delta^{\prime}\delta})\propto(s-S)^{2}$, and ${\cal F}(x,\alpha,\beta)$ is defined in Eq. (3.12). It can be noticed from Eq. (3.2) that
$x\rightarrow\sqrt{\frac{\Gamma_{\delta\delta}}{2\det{\bf\Sigma}_{(4)}}}$ $\left[\left(\frac{D^{\prime}}{2}-D^{2}\right)\frac{B_{s}}{2\Gamma_{\delta\delta}}\right.-$ (C.6)
@@
where $\det{\bf\Sigma}_{(4)}$ is the $4^{\rm th}$-order coefficient of the Taylor expansion. The above equation is a finite quantity, while $\alpha\propto(s-S)^{2}$ and $\beta\propto(s-S)^{-1}$. It follows that in the limit $s\rightarrow S$, the exponential term appearing in the integral of Eq. (3.12) behaves as a Gaussian infinitely wide and centered at $\infty$- ∞. Note, however, that the width of this Gaussian, i.e. $\sqrt{2/\alpha}$, and the center position, i.e. $-\beta$, compensate for each other,
$\alpha\beta^{2}=S\frac{\Gamma_{\delta^{\prime}\delta^{\prime}}}{\Gamma_{\delta^{\prime}\delta}^{2}}\left(\frac{B_{S}}{2S}-B^{\prime}\right)^{2}\rightarrow\frac{S}{\Gamma_{\delta\delta}}\left(\frac{B_{S}}{2S}-B^{\prime}\right)^{2}.$ (C.7)
This means that the exponential contributes to the integral as a finite positive quantity in the entire range from $\infty$- ∞ to $x$. As a consequence, the upper integration limit, $x$, is negligible and can be considered $x\rightarrow 0$. The same argument applies to the error function: the scale range in which it differs from -1 is negligible, therefore we can consider ${\rm erf}(t)\rightarrow-1$ over all the scales of interest. In this limit, the integral of Eq. (3.12) becomes
${\cal F}(x,\alpha,\beta)$ $\rightarrow-\int_{-\infty}^{0}{\rm d}t\,t\,e^{-\alpha(t+\beta)^{2}}$ (C.8)
@@
from which, the limit for $s\rightarrow S$ of Eq. (C.5), is
$I_{3}\rightarrow\frac{\sqrt{\Gamma_{\delta\delta}}}{4\pi S}e^{-\alpha\beta^{2}}+\frac{{\rm erf}(\sqrt{\alpha}\beta)+1}{4\sqrt{s\pi S}}\left(\frac{B_{S}}{2S}-B^{\prime}\right)\,.$ (C.9)
Substituting the four $I_{i}$ terms computed here in Eq. (C.2), we find that Eq. (3.2), in the limit $s\rightarrow S$, becomes
$\lim_{s\rightarrow S}{\cal P}(s)=e^{-B^{2}_{S}/2S}$ $\left\llbracket\frac{\sqrt{\Gamma_{\delta\delta}}}{2\pi S}\exp\left[-\frac{S}{2\Gamma_{\delta\delta}}\left(\frac{B_{S}}{2S}-B^{\prime}_{S}\right)^{2}\right]\right.+$ (C.10)
@@
$\frac{DB_{S}^{2}+SB^{\prime}-B_{S}B}{2\Gamma_{\delta\delta}}=\frac{S}{2\Gamma_{\delta\delta}}\left(\frac{B_{S}}{2S}-B^{\prime}_{S}\right)^{2}+\frac{B_{S}}{2S}.$ (C.11)
C.2 $s\rightarrow 0$ limit
We now explore the $s\rightarrow 0$ limit of ${\cal P}(s)$, Eq. (3.2). Note that $S>s$ by construction, so $R_{S}<R_{s}$. Considering $S$ as a finite quantity, $R_{S}\ll R_{s}\rightarrow\infty$ as $s\rightarrow 0$, entailing that
$C$ $=\int\frac{{\rm d}k}{2\pi^{2}}k^{2}P(k)W(kR_{S})W^{*}(kR_{s})$ (C.12)
$\rightarrow\int\frac{{\rm d}k}{2\pi^{2}}k^{2}P(k)W(kR_{s})={\cal O}(s)\,,$ (C.13)